OpenStudy (anonymous):

how do i factor Z^2+3i

OpenStudy (anonymous):

over what? complex numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lorda mercy you have to find \(\sqrt{3i}\) then right?

OpenStudy (anonymous):

yes lol exactly

OpenStudy (anonymous):

doesn't i^2 result in a negative number though?

OpenStudy (anonymous):

yes i^2 =-1...but I have to factor 3i, not square it

OpenStudy (anonymous):

well by the looks of things you have to find the square root of negative 3i

OpenStudy (anonymous):

to find z

OpenStudy (anonymous):

actually even that is not right is it? if it was \[z^2-3i\] then you would factor as \(z+\sqrt{3i})(z-\sqrt{3i})\)

OpenStudy (anonymous):

you actually need \(\sqrt{-3i}\)

OpenStudy (anonymous):

which is not really any different from finding \(\sqrt{3i}\)

OpenStudy (anonymous):

your right sat but 3i is positive.....(z^2+3i)

OpenStudy (anonymous):

lol no \(3i\) is not positive

OpenStudy (anonymous):



OpenStudy (anonymous):

write in trig form or exponential form as \[3\cos(\frac{3\pi}{2}+i\sin(\frac{3\pi}{2}))\] then take the real square root of 3 and divide the angle by 2

OpenStudy (anonymous):

one of the answers will be \[\sqrt{3}\left (\cos(\frac{3\pi}{4})+i\sin(\frac{3\pi}{4})\right)\]

OpenStudy (anonymous):

the other will be half way across the circle

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

I think I follow, but it somewhat leads me back to my original statement about the factors being positive

OpenStudy (anonymous):

complex numbers are not positive or negative

OpenStudy (anonymous):

suppose you wanted to factor \(z^2+4\) you could do this by writing \[z^2+4=0,z^2=-4,z=\pm\sqrt{-4}=\pm2i\] and so \[z^2+4=(z+2i)(z-2i)\]

OpenStudy (anonymous):

because \(-4\) has two complex roots, \(2i\) and \(-2i\) you have to do the same thing here, except instead of 4 you have \(3i\)

OpenStudy (anonymous):

meaning you want the two complex roots of \(-3i\)

OpenStudy (anonymous):

got you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes and thank you so much.

OpenStudy (anonymous):

yw